Evaluate the expressions. Round to the nearest hundredth.

log5^2
log2^5

Answer by Alan3354(21583)

Evaluate the expressions. Round to the nearest hundredth.
log5^2
= log(25)
= 1.40 (via calculator)
------------
log2^5
= log(32)
= 1.505....